OpenStudy (anonymous):

Can someone please help me!!!! :)) Change 4x - 3y = -15 to slope-intercept form and state the slope and y-intercept. You must show all work

OpenStudy (kropot72):

The slope intercept form is y = mx + b where m is the slope and b is the y-intercept. 4x - 3y = -15 First subtract 4x from both sides of the equation, as follows; 4x -4x - 3y = -15 - 4x This simplifies to -3y = -4x - 15 This last equation can be turned into the required form by dividing both sides by -3:\[\frac{-3y}{-3}=\frac{-4x-15}{-3}\] Can you simplify the last equation?

OpenStudy (anonymous):

Thank you!!!!!!!!!!!!!!!!!!!!!!! and no

OpenStudy (kropot72):

Let's take it in steps, starting with the fraction -3y/-3. What do you get if you multiply the numerator and the denominator by -1.